# ReliableMath: Benchmark of Reliable Mathematical Reasoning on Large Language Models

### v1

@@ F.1 Solvable Data Collection
  • AIME 2024 [1]  [^3] https://huggingface.co/datasets/AI-MO/aimo-validation-aime serves as a benchmark comprising exceptionally challenging mathematical problems from the American Invitational Mathematics Examination (AIME). These problems are specifically designed to evaluate advanced problem-solving abilities and are considerably more difficult than standard high school mathematics questions. Solving these items requires sophisticated reasoning and the application of advanced strategies.
@@
  • AMC dataset [2]  [^4] https://huggingface.co/datasets/AI-MO/aimo-validation-amc serves as an internal validation set during participation in the AIMO progress prize competition. The dataset contains 83 problems from AMC12 2022 and AMC12 2023 extracted from the AOPS wiki page  [^5] https://artofproblemsolving.com/wiki/index.php/AMC_12_Problems_and_Solutions.
@@
  • MinervaMath [22]  [^6] https://huggingface.co/datasets/math-ai/minervamath includes 272 college-level problems in physical, engineering, chemistry, economics, and other sciences that require quantitative reasoning.
@@
  • MATH [16]  [^7] https://github.com/hendrycks/math/ is a challenging mathematical dataset with competitive mathematics problems, consisting of 7,500 training samples and 5,000 test samples. Each problem in MATH also has a full step-by-step solution which can be used to teach models to generate answer derivations and explanations across several subjects, including algebra, geometry, number theory, counting and probability, calculus, etc.
F.2 Data Format
@@
Appendix G Human Evaluation Instruction
@@ G.1 Description of Task
Then, the original question, ground-truth answer, and the auxiliary information, including the rewriting process and the modified positions, will be presented to the evaluator to assist in the manual annotation process of verifying whether the question is unsolvable, thereby improving the quality of human evaluation. The detailed format of data presented to the human annotator will be explained in Section G.2.
The evaluator’s goal is to combine the above-mentioned information to determine whether the constructed unsolvable problem is indeed unsolvable. The detailed evaluation principles will be explained in Section G.3.
G.2 Evaluation Data Format and Explanation
@@ G.3 Principle of Evaluation
Appendix H Experiments
H.1 Model Details
We attach the model links of employed LLMs as follows:
H.2 Completed Experimental Results
The completed experiments on respective ReliableMath subsets of AIME, AMC, MATH, and Minerva are presented in Table 6, 7, 8, 9, and Fig. 12.
Figure 12: Illustrations of lengths of generations (columns (a)), reliability figures of Precision (column (b)) and Prudence (column (c)) on various ReliableMath subsets. In column (a), slv and usl denote solvable and unsolvable problems, while std and real denote using standard and reliable prompts, respectively. Only four representative LLMs are presented. x-axis and y-axis in reliability figures denote Prec.($\mathcal{A}$)/Prud.($\mathcal{A}$) and Prec.($\mathcal{U}$)/Prud.($\mathcal{U}$) as in Eq. 1 / 2. We employ different colors to indicate different LLMs. ◆ and ★ in figures represent using standard and reliable prompts. The dashed lines connecting two points in the same color represent the trajectory of reliability changes from the standard prompt ◆ to the reliable prompt ★ of an LLM.
Table 6: Reliability evaluations on ReliableMath AIME Solve (SLV) and Unsolve (USV) subsets.
LLMs Prompt AIME SLV ($\mathcal{A}$) AIME USV ($\mathcal{U}$) Prec. Prud.
$\mathcal{S}$$\mathcal{R}$/$\mathcal{U}$ Len. Prec. Prud. $\mathcal{S}$$\mathcal{R}$/$\mathcal{U}$ Len. Prec. Prud.
DeepSeek-R1 standard 25/0/5 7591 0.833 0.000 0/0/132 11070 0.000 0.000 0.417 0.000
reliable 21/0/9 7569 0.700 0.000 52/3/77 8527 0.394 0.023 0.547 0.011
o3-mini standard 19/0/11 3456 0.633 0.000 0/0/132 8467 0.000 0.000 0.317 0.000
reliable 22/0/8 3890 0.733 0.000 33/5/94 5214 0.250 0.037 0.492 0.019
Distill-32B standard 20/0/10 8341 0.666 0.000 0/0/132 20630 0.000 0.000 0.333 0.000
reliable 19/0/11 12569 0.633 0.000 34/2/96 15676 0.258 0.015 0.445 0.008
Distill-14B standard 18/0/12 12829 0.600 0.000 0/0/132 26297 0.000 0.000 0.300 0.000
reliable 16/0/14 15456 0.533 0.000 46/1/85 20259 0.348 0.008 0.441 0.004
Distill-7B standard 14/0/16 11522 0.466 0.000 0/0/132 11131 0.000 0.000 0.233 0.000
reliable 11/0/19 12001 0.366 0.000 0/0/132 11798 0.000 0.000 0.183 0.000
Distill-1.5B standard 8/0/22 13564 0.266 0.000 0/0/132 13359 0.000 0.000 0.133 0.000
reliable 6/0/24 13877 0.200 0.000 0/0/132 13542 0.000 0.000 0.100 0.000
DeepSeek-V3 standard 0/0/132 3506 0.466 0.000 0/0/132 3506 0.000 0.000 0.233 0.000
reliable 14/0/16 3009 0.466 0.000 24/0/108 3005 0.181 0.000 0.324 0.000
GPT-4o standard 2/0/28 765 0.066 0.000 0/0/132 741 0.000 0.000 0.033 0.000
reliable 3/0/27 689 0.100 0.000 22/5/105 933 0.166 0.037 0.133 0.019
Qwen2.5-7B standard 9/0/21 1278 0.300 0.000 0/0/132 1234 0.000 0.000 0.150 0.000
reliable 5/0/25 1494 0.166 0.000 1/0/131 1458 0.008 0.000 0.087 0.000
Qwen2.5-1.5B standard 2/0/28 892 0.066 0.000 0/0/132 949 0.000 0.000 0.033 0.000
reliable 3/0/27 953 0.100 0.000 0/0/132 887 0.000 0.000 0.050 0.000
Table 7: Reliability evaluations on ReliableMath AMC Solve (SLV) and Unsolve (USV) subsets.
LLMs Prompt AMC SLV ($\mathcal{A}$) AMC USV ($\mathcal{U}$) Prec. Prud.
$\mathcal{S}$$\mathcal{R}$/$\mathcal{U}$ Len. Prec. Prud. $\mathcal{S}$$\mathcal{R}$/$\mathcal{U}$ Len. Prec. Prud.
DeepSeek-R1 standard 77/0/6 5228 0.927 0.000 0/0/295 7920 0.000 0.000 0.464 0.000
reliable 75/0/8 4901 0.903 0.000 152/2/141 5339 0.515 0.007 0.709 0.003
o3-mini standard 71/0/12 2224 0.855 0.000 1/0/294 5559 0.003 0.000 0.429 0.000
reliable 73/0/10 2908 0.879 0.000 90/1/204 5673 0.305 0.003 0.592 0.002
Distill-32B standard 71/0/12 5814 0.855 0.000 0/0/295 17691 0.000 0.000 0.428 0.000
reliable 68/0/15 5654 0.819 0.000 107/0/188 13137 0.363 0.000 0.591 0.000
Distill-14B standard 71/0/12 7376 0.855 0.000 0/0/295 20426 0.000 0.000 0.428 0.000
reliable 67/0/16 7737 0.807 0.000 134/0/161 14516 0.454 0.000 0.631 0.000
Distill-7B standard 67/0/16 6924 0.807 0.000 0/0/295 7076 0.000 0.000 0.404 0.000
reliable 53/0/30 7923 0.638 0.000 0/0/295 7993 0.000 0.000 0.319 0.000
Distill-1.5B standard 38/0/45 10180 0.457 0.000 0/0/295 10142 0.000 0.000 0.229 0.000
reliable 38/0/45 10123 0.457 0.000 0/0/295 10035 0.000 0.000 0.229 0.000
DeepSeek-V3 standard 67/0/16 2174 0.807 0.000 0/0/295 2575 0.000 0.000 0.404 0.000
reliable 69/0/14 1998 0.831 0.000 86/2/207 2189 0.291 0.007 0.561 0.003
GPT-4o standard 38/0/45 1146 0.457 0.000 0/0/295 1009 0.000 0.000 0.229 0.000
reliable 37/0/46 743 0.445 0.000 89/10/196 895 0.301 0.033 0.374 0.017
Qwen2.5-7B standard 41/0/42 1079 0.494 0.000 0/0/295 1058 0.000 0.000 0.247 0.000
reliable 48/0/35 919 0.578 0.000 5/0/290 935 0.017 0.000 0.298 0.000
Qwen2.5-1.5B standard 45/0/38 815 0.542 0.000 0/0/295 840 0.000 0.000 0.271 0.000
reliable 39/0/44 827 0.469 0.000 0/0/295 854 0.000 0.000 0.235 0.000
Table 8: Reliability evaluations on ReliableMath MATH Solve (SLV) and Unsolve (USV) subsets.
LLMs Prompt MATH SLV ($\mathcal{A}$) MATH USV ($\mathcal{U}$) Prec. Prud.
$\mathcal{S}$$\mathcal{R}$/$\mathcal{U}$ Len. Prec. Prud. $\mathcal{S}$$\mathcal{R}$/$\mathcal{U}$ Len. Prec. Prud.
DeepSeek-R1 standard 93/0/7 3116 0.930 0.000 0/0/318 5762 0.000 0.000 0.465 0.000
reliable 93/0/7 2709 0.930 0.000 189/3/126 3789 0.594 0.009 0.762 0.005
o3-mini standard 89/0/11 1247 0.890 0.000 0/0/318 6184 0.000 0.000 0.445 0.000
reliable 91/0/9 817 0.910 0.000 104/0/214 5219 0.327 0.000 0.619 0.000
Distill-32B standard 90/0/10 4313 0.900 0.000 0/0/318 12126 0.000 0.000 0.450 0.000
reliable 87/0/13 4523 0.870 0.000 157/0/161 7813 0.494 0.000 0.682 0.000
Distill-14B standard 84/0/16 4831 0.840 0.000 0/0/318 14465 0.000 0.000 0.420 0.000
reliable 82/0/18 5150 0.820 0.000 166/0/152 9349 0.522 0.000 0.671 0.000
Distill-7B standard 85/0/15 4082 0.850 0.000 0/0/318 4536 0.000 0.000 0.425 0.000
reliable 84/0/16 4427 0.840 0.000 0/0/318 4893 0.000 0.000 0.420 0.000
Distill-1.5B standard 67/0/33 6500 0.670 0.000 0/0/318 7425 0.000 0.000 0.335 0.000
reliable 64/0/36 7135 0.640 0.000 0/0/318 7769 0.000 0.000 0.320 0.000
DeepSeek-V3 standard 89/0/11 1147 0.890 0.000 0/0/318 1795 0.00 0.000 0.445 0.000
reliable 90/0/10 1156 0.900 0.000 135/0/183 1410 0.424 0.000 0.662 0.000
GPT-4o standard 72/0/28 552 0.720 0.000 0/0/318 736 0.000 0.000 0.360 0.000
reliable 70/1/29 506 0.700 0.010 120/6/192 490 0.377 0.019 0.539 0.014
Qwen2.5-7B standard 79/0/21 674 0.790 0.000 0/0/318 695 0.000 0.000 0.395 0.000
reliable 80/0/20 690 0.800 0.000 2/0/316 773 0.006 0.000 0.403 0.000
Qwen2.5-1.5B standard 68/0/32 577 0.680 0.000 0/0/318 649 0.000 0.000 0.340 0.000
reliable 68/0/32 593 0.680 0.000 2/0/316 613 0.006 0.000 0.343 0.000
Table 9: Reliability evaluations on ReliableMath Minerva Solve (SLV) and Unsolve (USV) subsets.
LLMs Prompt Minerva SLV ($\mathcal{A}$) Minerva USV ($\mathcal{U}$) Prec. Prud.
$\mathcal{S}$$\mathcal{R}$/$\mathcal{U}$ Len. Prec. Prud. $\mathcal{S}$$\mathcal{R}$/$\mathcal{U}$ Len. Prec. Prud.
DeepSeek-R1 standard 39/0/61 3117 0.390 0.000 0/0/357 4271 0.000 0.000 0.195 0.000
reliable 41/0/59 2872 0.410 0.000 212/0/145 2575 0.594 0.000 0.502 0.000
o3-mini standard 29/0/71 591 0.290 0.000 0/0/357 1902 0.000 0.000 0.145 0.000
reliable 38/2/60 530 0.380 0.020 96/0/261 1918 0.269 0.000 0.324 0.010
Distill-32B standard 42/0/58 3971 0.420 0.000 0/0/357 11819 0.000 0.000 0.210 0.000
reliable 40/0/60 2808 0.400 0.000 163/0/194 5259 0.457 0.000 0.428 0.000
Distill-14B standard 37/0/63 5571 0.370 0.000 0/0/357 12923 0.000 0.000 0.185 0.000
reliable 32/0/68 3305 0.320 0.000 166/0/191 6025 0.465 0.000 0.393 0.000
Distill-7B standard 25/0/75 5997 0.250 0.000 0/0/357 6101 0.000 0.000 0.125 0.000
reliable 32/0/68 4941 0.320 0.000 3/0/354 5036 0.008 0.000 0.164 0.000
Distill-1.5B standard 18/0/82 9155 0.180 0.000 0/0/357 9616 0.000 0.000 0.090 0.000
reliable 16/0/84 9627 0.160 0.000 0/0/357 9727 0.000 0.000 0.080 0.000
DeepSeek-V3 standard 35/0/65 582 0.350 0.000 1/0/356 828 0.003 0.000 0.176 0.000
reliable 35/0/65 488 0.350 0.000 171/1/185 564 0.479 0.003 0.414 0.001
GPT-4o standard 32/0/68 510 0.320 0.000 1/0/356 658 0.003 0.000 0.161 0.000
reliable 34/1/65 480 0.340 0.010 138/6/213 467 0.386 0.017 0.363 0.013
Qwen2.5-7B standard 27/0/73 651 0.270 0.000 0/0/357 689 0.000 0.000 0.135 0.000
reliable 25/0/75 673 0.250 0.000 22/0/335 695 0.061 0.000 0.156 0.000
Qwen2.5-1.5B standard 26/0/74 694 0.260 0.000 0/0/357 722 0.000 0.000 0.130 0.000
reliable 22/0/78 743 0.220 0.000 14/0/343 749 0.039 0.000 0.130 0.000
Figure 13: Illustrations of (a) precision of unsolvable problems constructed by two rewriting types in Sec. 3.2 (removal & contradiction), and (b) precision of unsolvable problems of two difficulty levels rated by human experts in Sec. 3.3 (0: simple & 1: hard) on all LLMs employing reliable prompts.
H.3 Training Settings
During training, ADAM parameter update is used in a mini-batch mode. The initial learning rate of 1e-4 is utilized with the 0.05 warm-up ratio and 0.01 weight decay of the ADAM optimizer. When training the models, we set epochs to 3 to ensure that all the models can be trained to convergence. All LLMs are quantified using brain floating point (bf16) to load and save parameters. The vLLM library [21]  [^16] https://github.com/vllm-project/vllm is utilized to accelerate the generation.
@@ H.3 Training Settings
AIME [1] 30 30 85 85 170 70 71 141 67 65 132
AMC [2] 83 83 241 241 482 185 192 377 131 164 295
MATH500 [16] 500 100 273 273 546 233 216 449 154 164 318
Minerva [31] 272 100 254 254 508 207 201 408 185 172 357
@@
AIME [1] 67 (24/43) 65 (22/43) 132 (46/86)
AMC [2] 131 (47/84) 164 (60/104) 295 (107/188)
MATH500 [16] 154 (77/77) 164 (75/89) 318 (152/166)
Minerva [31] 185 (87/96) 172 (76/96) 357 (163/192)
@@
Dataset ${\mathcal{Q}_{a}}^{(t)}$ Step 1 Step 2 ${\mathcal{Q}_{u}}^{(t)}$
AIME [1] 975 1950 961
AMC [2] 3264 6528 3185
MATH [16] 2298 4596 961
Total 6537 13074 5085

### v2

@@ 9.1 Solvable Data Collection
  • AIME 2024 (aimo2024aime)  [^4] https://huggingface.co/datasets/AI-MO/aimo-validation-aime serves as a benchmark comprising exceptionally challenging mathematical problems from the American Invitational Mathematics Examination (AIME). These problems are specifically designed to evaluate advanced problem-solving abilities and are considerably more difficult than standard high school mathematics questions. Solving these items requires sophisticated reasoning and the application of advanced strategies.
@@
  • AMC dataset (aimo2024amc)  [^5] https://huggingface.co/datasets/AI-MO/aimo-validation-amc serves as an internal validation set during participation in the AIMO progress prize competition. The dataset contains 83 problems from AMC12 2022 and AMC12 2023 extracted from the AOPS wiki page  [^6] https://artofproblemsolving.com/wiki/index.php/ AMC_12_Problems_and_Solutions.
@@
  • MinervaMath (lewkowycz2022solving)  [^7] https://huggingface.co/datasets/math-ai/minervamath includes 272 college-level problems in physical, engineering, chemistry, economics, and other sciences that require quantitative reasoning.
@@
  • MATH (hendrycks2021measuring)  [^8] https://github.com/hendrycks/math/ is a challenging mathematical dataset with competitive mathematics problems, consisting of 7,500 training samples and 5,000 test samples. Each problem in MATH also has a full step-by-step solution which can be used to teach models to generate answer derivations and explanations across several subjects, including algebra, geometry, number theory, counting and probability, calculus, etc.
9.2 Data Format
@@
10 Ethics Statement
11 Potential Social Impact
12 Human Evaluation Instruction
@@ 12.1 Description of Task
Then, the original question, ground-truth answer, and the auxiliary information, including the rewriting process and the modified positions, will be presented to the evaluator to assist in the manual annotation process of verifying whether the question is unsolvable, thereby improving the quality of human evaluation. The detailed format of data presented to the human annotator will be explained in Section 12.2.
The evaluator’s goal is to combine the above-mentioned information to determine whether the constructed unsolvable problem is indeed unsolvable. The detailed evaluation principles will be explained in Section 12.3.
12.2 Evaluation Data Format and Explanation
@@ 12.3 Principle of Evaluation
AIME 30 30 85 85 170 70 71 141 67 65 132
AMC 83 83 241 241 482 185 192 377 131 164 295
MATH500 500 100 273 273 546 233 216 449 154 164 318
Minerva 272 100 254 254 508 207 201 408 185 172 357
@@
AIME 67 (24/43) 65 (22/43) 132 (46/86)
AMC 131 (47/84) 164 (60/104) 295 (107/188)
MATH500 154 (77/77) 164 (75/89) 318 (152/166)
Minerva 185 (87/96) 172 (76/96) 357 (163/192)
@@
13 Experiments
13.1 Model Details
We attach the model links of employed LLMs as follows:
13.2 Completed Experimental Results
The completed experiments on respective ReliableMath subsets of AIME, AMC, MATH, and Minerva are presented in Table 11, 12, 13, 14, and Fig. 12.
13.3 Training Settings
All the LLMs are trained using Llama-Factory (zheng2024llamafactory).  [^17] https://github.com/hiyouga/LLaMA-Factory During training, ADAM parameter update is used in a mini-batch mode with batch size = 16. The initial learning rate of 1e-5 is utilized with the 0.05 warm-up ratio and 0.01 weight decay of the ADAM optimizer. When training the models, we set epochs to 3 to ensure that all the models can be trained to convergence. All LLMs are quantified using brain floating point (bf16) to load and save parameters. The vLLM library (kwon2023efficient)  [^18] https://github.com/vllm-project/vllm is utilized to accelerate the generation.
Dataset $\mathcal{Q}_{a}$} Step 1 Step 2 $\mathcal{Q}_{u}$}
AIME 975 1950 961
AMC 3264 6528 3185
MATH 2298 4596 961
Total 6537 13074 5085